OpenStudy (anonymous):

int_{0.1}^{0.7}cosx ln(2sinx)dx

OpenStudy (anonymous):

\[\int\limits_{0.1}^{0.7}cosx \ln(2sinx)dx\]

OpenStudy (anonymous):

Let \(u=2\sin x\), so \(\dfrac{1}{2}~du=\cos x~dx\).

OpenStudy (anonymous):

Integral changes to \[\frac{1}{2}\int_{2\sin(0.1)}^{2\sin(0.7)}\ln u~du\] Then integrate by parts, or use the formula for \(\int\ln u~du\) if you know it.

OpenStudy (anonymous):

could you please explain more

OpenStudy (anonymous):

As in continue? or explain what I did so far?

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

\[\frac{1}{2}\int_{2\sin(0.1)}^{2\sin(0.7)}\ln u~du=\frac{1}{2}\left[fg-\int g~df\right]_{2\sin(0.1)}^{2\sin(0.7)}\] where you have \[\begin{matrix}f=\ln u&&&dg=du\\df=\frac{du}{u}&&&g=u\end{matrix}\] \[\frac{1}{2}\left[fg-\int g~df\right]_{2\sin(0.1)}^{2\sin(0.7)}=\frac{1}{2}\left[\left[u\ln u\right]_{2\sin(0.1)}^{2\sin(0.7)}-\int_{2\sin(0.1)}^{2\sin(0.7)}u\left(\frac{du}{u}\right)\right]\] Simplified a bit, you have \[\frac{1}{2}\left[\left[u\ln u\right]_{2\sin(0.1)}^{2\sin(0.7)}-\int_{2\sin(0.1)}^{2\sin(0.7)}du\right]\\ \frac{1}{2}\Bigg[\left[u\ln u\right]_{2\sin(0.1)}^{2\sin(0.7)}-\left[u\right]_{2\sin(0.1)}^{2\sin(0.7)}\Bigg]\]

OpenStudy (anonymous):

so what you get for answer

OpenStudy (anonymous):

just checking